OpenStudy (anonymous):

find an equation of the line containging these points (-2,-4) and (-9,-6) need be in slope-intercept form

OpenStudy (saifoo.khan):

\[m=\frac{-6+4}{-9+2}\]\[m=\frac{-2}{-7}\]\[m = \frac{2}{7}\]

OpenStudy (mimi_x3):

m = y^2 - y1 / x^2-x^1 = 2/7 Find then y = 2/7x -24/7

OpenStudy (mimi_x3):

am i right ? xD

OpenStudy (saifoo.khan):

u are Late. Congrats yes.

OpenStudy (mimi_x3):

Excuse me l8 , i was finding my calculator and it takes time xDD

OpenStudy (saifoo.khan):

LOL, why & where do u hide your calc.

OpenStudy (mimi_x3):

LOLs it was in my bag but I forgot it was in it so i was finding it xD

OpenStudy (saifoo.khan):

Dumb.

OpenStudy (anonymous):

dw abt it mimi saifooo is good at stealing the #1 spot he's had it for a long time now, he's mastered the art of calculator and posting quickly